# A Monocular Wide-FoV SLAM System Based on Point Uncertainty and Pose Uncertainty

### v1

@@ VI-A Evaluation on PALVIO Dataset
The PALVIO dataset consists of $10$ indoor sequences, recorded using two hand-held PAL cameras (FoV: $360^{\circ}{\times}(40^{\circ}{\sim}120^{\circ}$, resolution: $1280{\times}960$) and an IMU sensor ($200Hz$). With the data captured by the motion capture system as the ground truth, we compare our P2U-SLAM system with four open-source state-of-the-art vSLAM/VIO methods SVO2.0 [3], VINS-Mono [9], ORB-SLAM3 (monocular mode) [23], and LF-VIO [26]. Absolute Translation Error (ATE) and Relative Pose Error (RPE) are used as the evaluation metrics for assessment.
@@
As shown in Table I, P2U-SLAM has the lowest average error in ATE and RPE metrics. In detail, P2U-SLAM reaches the best precision on all sequences in ATE and RPEt metrics, and most sequences in RPEr metrics. On a handful of sequences where P2U-SLAM ranks 2nd in RPEr metrics, the SLAM/VIO methods (SVO2.0 and LF-VIO) performing best are all fusing IMU measurements, while P2U-SLAM reaches comparable precision with the only monocular visual PAL input. Furthermore, ID02, ID04, and ID06 are chosen as representative sequences for exhibition in Fig. 5.
@@
Apart from SVO2.0 whose data are cited from [26], all other SLAM/VIO frameworks are tested $10$ times on the PALVIO dataset and the outcome of experiments is expressed as colored squares in Fig. 6 to evaluate the robustness of these frameworks. The fewer turns of failure are, the more robust the corresponding system is. P2U-SLAM and LF-VIO are naturally able to process map points from the negative half-plane to fully utilize the whole FoV of the PAL image, due to the adaptation to the Taylor camera model. The larger FoV makes P2U-SLAM and LF-VIO stably track in rapid rotation and motion. Therefore, P2U-SLAM and LF-VIO both succeed in every trial of experiments. Meanwhile, with the help of point and pose uncertainty, the P2U-SLAM can suppress the influence on optimization from outliers. P2U-SLAM achieves the lowest variance in ten times testing. In summary, P2U-SLAM is verified to be robust to wide-FoV PAL input and notably stands out among the compared methods on this dataset.
VI-B Evaluation on TUM-VI Benchmark
Three types of sequences of room, corridor, and magistrale are selected from the TUM-VI dataset to evaluate the performance of P2U-SLAM in various complicated environments such as rapid motion, duplicate textures, scene switches, and brightness changes. The resolution and FoV of fisheye image input are respectively set $512{\times}512$ and $360^{\circ}{\times}(0^{\circ}{\sim}95^{\circ})$, and the frequency of IMU data is $200Hz$ on the TUM-VI benchmark. The ground truth captured by the motion capture system is available for complete trajectory on room sequences, while unavailable except at the starting and ending points on corridor and magistrale sequences. Therefore, ATE serves as the suitable metric to measure the cumulative drift along the whole trajectory.
TABLE II: Performance comparison on the public TUM-VI benchmark [60] (RMS ATE in $m$).
Sequences VINS- LF- ORB- P2U- Length
Mono VIO SLAM31 SLAM (m)
[9] [26] [23] (ours)
room1 0.089 0.094 0.074 0.071 146
room2 0.107 0.129 0.048 0.047 142
room3 0.129 0.114 0.060 0.049 135
room4 0.251 0.361 0.078 0.079 68
room5 0.219 0.220 0.072 0.075 131
room6 0.097 0.120 0.073 0.074 67
corridor1 0.771 0.380 0.358 0.065 305
corridor2 0.872 0.413 0.073 0.065 322
corridor3 0.635 0.585 0.269 0.070 300
corridor4 0.186 0.471 0.231 0.069 114
corridor5 0.845 0.484 0.085 0.060 270
magistrale1 18.68 4.805 1.933 0.085 918
magistrale2 1.956 0.543 0.748 0.086 561
magistrale3 0.756 1.894 0.969 0.086 566
magistrale4 1.025 0.780 0.502 0.084 688
magistrale5 1.690 1.135 1.331 0.081 458
magistrale6 4.032 4.167 0.176 0.093 711
@@ VI-B Evaluation on TUM-VI Benchmark
The RMS ATE results obtained are compared with the most relevant open-source systems in Table II. On room sequences, P2U-SLAM achieves comparable precision which is less than $0.1m$ with the state-of-the-art framework ORB-SLAM3 (monocular) [23]. However, the cumulative drift of the ORB-SLAM3 increases significantly with the length of the test sequence on corridor and magistrale sequences like other VIO methods, while the P2U-SLAM always maintains the ATE of no more than $0.1m$ on every sequence. Considering that P2U-SLAM and ORB-SLAM3 use the same feature extraction method and multi-thread framework, the results in Fig. 7 illustrate that the uncertainty method adopted by P2U-SLAM achieves a high degree of error stability in trajectory tracking within one kilometer.
TABLE III: Ablation studies for point and pose uncertainty with metric ATE ($m$) on sequences corridor1$\sim$5 of TUM-VI benchmark [60].
@@
VI-C Ablation Study
TABLE IV: Running time of the main parts of P2U-SLAM compared to ORB-SLAM3 (mono) [23], measured on TUM-VI corridor2 [60] (mean time and standard deviation in $ms$).
Module P2U-SLAM ORB-SLAM3  [23]
Tracking 20.68$\pm$7.84 16.97$\pm$4.77
Local mapping 198.13$\pm$129.87 174.36$\pm$102.66
Place recognition 5.25$\pm$2.83 6.12$\pm$4.08
Loop Fusion 42.82$\pm$25.56 43.61$\pm$27.20
Loop Global BA 451.06$\pm$189.24 447.25$\pm$186.82
Uncertainty 0.83$\pm$0.55 /
@@ VI-C Ablation Study
VI-D Computing Time Analysis
In this subsection, we report the efficiency of P2U-SLAM. Table IV compares the running time of the main operations performed in P2U-SLAM with ORB-SLAM3, indicating that P2U-SLAM owns the ability to work in real-time at about $40{\sim}60$ frames and $3{\sim}6$ keyframes per second. Nevertheless, the average running time of P2U-SLAM is slightly larger than ORB-SLAM3 (mono) with around $22\%$ and $14\%$ in respectively tracking and local mapping module, since the introduction of uncertainty module in optimization of P2U-SLAM. This efficient performance highlights P2U-SLAM’s capability to maintain high accuracy with comparable speed, making it ideal for real-time applications.
@@ VII Conclusion
In this study, we propose P2U-SLAM, a framework for wide-FoV cameras that makes use of point and pose uncertainty to overcome weak feature correspondence introduced by wide overlap area. The proposed system establishes a complete visual SLAM framework, which includes initialization, tracking, local mapping, loop closing, and map merging modules. To handle 3D points from the negative place ($z{<}0$), the system inherits the Taylor camera model used in previous work. In addition, this work shows that it is mathematically reasonable and necessary to incorporate point and pose uncertainty during optimization. Moreover, the calculation and use of point and pose uncertainty are integrated into the SLAM framework, allowing P2U-SLAM to inhibit positioning accuracy and robustness that exceeds the state-of-the-art methods on two public wide-FoV SLAM datasets. There is still a lot of room for improvement in P2U-SLAM. Despite a significant decrease in cumulative drift in the kilometer-level series, P2U-SLAM is still a pure visual SLAM framework. It is challenging for P2U-SLAM to work in some scenarios such as elevators and subways. Besides, the visualization of the map generated by P2U-SLAM is sparse and limited in downstream applications compared to dense maps.
In the future, we intend to adapt P2U-SLAM to the visual-inertial framework to enhance the robustness of our algorithm in more extreme scenarios where vision fails. In addition, the uncertainty method of P2U-SLAM is strongly related to the current hot research on 3D Gaussian sputtering, so achieving real-time dense mapping with higher efficiency from P2U-SLAM is also one of our exploration directions.

### v2

@@ VI-A Evaluation on PALVIO Dataset
The PALVIO dataset consists of $10$ indoor sequences, recorded using two hand-held PAL cameras (FoV: $360^{\circ}{\times}(40^{\circ}{\sim}120^{\circ}$, resolution: $1280{\times}960$)) and an IMU sensor ($200Hz$). With the data captured by the motion capture system as the ground truth, we compare our P2U-SLAM system with four open-source state-of-the-art vSLAM/VIO methods SVO2.0 [forster2016svo], VINS-Mono [qin2018vins], ORB-SLAM3 (monocular mode) [campos2021orb], LF-VIO [wang2022lf], and LF-VISLAM [wang2023lfvislam]. Absolute Translation Error (ATE) and Relative Pose Error (RPE) are used as the evaluation metrics for assessment.
@@
As shown in Table I, P2U-SLAM has the lowest average error in ATE and RPE metrics. In detail, P2U-SLAM reaches the best precision on all sequences in ATE and RPEt metrics, and most sequences in RPEr metrics. On a handful of sequences where P2U-SLAM ranks 2nd in RPEr metrics, the SLAM/VIO methods (SVO2.0, LF-VIO, and LF-VISLAM) performing best are all fusing IMU measurements, while P2U-SLAM reaches comparable precision with the only monocular visual PAL input. Furthermore, ID02, ID04, and ID08 are chosen as representative sequences for exhibition in Fig. 5.
@@
Apart from SVO2.0 whose data are cited from [wang2022lf], all other SLAM/VIO frameworks are tested $10$ times on the PALVIO dataset and the outcome of experiments is expressed as colored squares in Fig. 6 to evaluate the robustness of these frameworks. The fewer turns of failure there are, the more robust the corresponding system is. P2U-SLAM, LF-VIO, and LF-VISLAM are naturally able to process map points from the negative half-plane to fully utilize the whole FoV of the PAL image, due to the adaptation to the Taylor camera model. The larger FoV has made a great contribution to stably tracking in rapid rotation and motion. Therefore, P2U-SLAM, LF-VIO, and LF-VISLAM succeed in every trial of experiments. However, the robustness performance of these three methods is not the same. With the help of point and pose uncertainty, the P2U-SLAM can suppress the influence on optimization from outliers or weak constraints, and P2U-SLAM finally achieves the lowest variance in ten times of testing.
In summary, P2U-SLAM has been verified to be robust to wide-FoV PAL input and notably stands out among the compared methods in this dataset with all methods under Sim3 alignment.
TABLE II: Performance comparison on the public TUM-VI benchmark [schubert2018tum] (RMS ATE in $m$).
Sequences VINS- LF- LF- ORB- P2U- Length
Mono VISLAM VIO SLAM31 SLAM (m)
[qin2018vins] [wang2023lfvislam] [wang2022lf] [campos2021orb] (ours)
room1 0.089 0.089 0.094 0.074 0.071 146
room2 0.107 0.117 0.129 0.048 0.047 142
room3 0.129 0.123 0.114 0.060 0.049 135
room4 0.251 0.059 0.361 0.078 0.079 68
room5 0.219 0.174 0.220 0.072 0.075 131
room6 0.097 0.113 0.120 0.073 0.074 67
corridor1 0.771 0.476 0.380 0.358 0.065 305
corridor2 0.872 0.569 0.413 0.073 0.065 322
corridor3 0.635 0.638 0.585 0.269 0.070 300
corridor4 0.186 0.520 0.471 0.231 0.069 114
corridor5 0.845 0.487 0.484 0.085 0.060 270
magistrale1 18.68 0.747 4.805 1.933 0.085 918
magistrale2 1.956 0.531 0.543 0.748 0.086 561
magistrale3 0.756 0.534 1.894 0.969 0.086 566
magistrale4 1.025 0.736 0.780 0.502 0.084 688
magistrale5 1.690 0.885 1.135 1.331 0.081 458
magistrale6 4.032 0.727 4.167 0.176 0.093 711
@@
VI-B Evaluation on TUM-VI Benchmark
Three types of sequences of room, corridor, and magistrale are selected from the TUM-VI dataset to evaluate the performance of P2U-SLAM in various complicated environments such as rapid motion, duplicate textures, scene switches, and brightness changes. The resolution and FoV of the fisheye image input are respectively set $512{\times}512$ and $360^{\circ}{\times}(0^{\circ}{\sim}95^{\circ})$, and the frequency of IMU data is $200Hz$ on the TUM-VI benchmark. The ground truth captured by the motion capture system is available for complete trajectory on room sequences, while unavailable except at the starting and ending points on corridor and magistrale sequences. Therefore, ATE serves as a suitable metric to measure the cumulative drift along the whole trajectory.
The RMS ATE results obtained are compared with the most relevant open-source systems in Table II. On room sequences, P2U-SLAM achieves comparable precision which is less than $0.1m$ with the state-of-the-art framework ORB-SLAM3 (monocular) [campos2021orb]. However, the cumulative drift of the ORB-SLAM3 increases significantly with the length of the test sequence on corridor and magistrale sequences like other VIO methods, while the P2U-SLAM always maintains the ATE of no more than $0.1m$ on every sequence. Considering that P2U-SLAM and ORB-SLAM3 use the same feature extraction method and multi-thread framework, the results in Fig. 7 illustrate that the uncertainty method adopted by P2U-SLAM achieves a high degree of error stability in trajectory tracking within one kilometer. It is also worth noting that, with the support of the complete loop closure function and IMU data, LF-VISLAM also maintains excellent consistency on the TUM-VI sequences with multi-scale changes. However, it is still slightly inferior compared to P2U-SLAM, which has the support of dual uncertainties.
@@ VI-B Evaluation on TUM-VI Benchmark
VI-C Ablation Study
[…]
@@ VI-C Ablation Study
VI-D Computing Time Analysis
In this subsection, we report the efficiency of P2U-SLAM. Table IV compares the running time of the main operations performed in P2U-SLAM with ORB-SLAM3, indicating that P2U-SLAM owns the ability to work in real-time at about $40{\sim}60$ frames and $3{\sim}6$ keyframes per second. Nevertheless, the average running time of P2U-SLAM is slightly larger than ORB-SLAM3 (mono) with around $22\%$ and $14\%$ in respectively the tracking and local mapping module, since the introduction of the uncertainty module in optimization of P2U-SLAM. This efficient performance highlights P2U-SLAM’s capability to maintain high accuracy with comparable speed, making it ideal for real-time applications.
TABLE IV: Running time of the main parts of P2U-SLAM compared to ORB-SLAM3 (mono) [campos2021orb], measured on TUM-VI corridor2 [schubert2018tum] (mean time and standard deviation in msms).
Module P2U-SLAM ORB-SLAM3  [campos2021orb]
[…]
@@ VII Conclusion
In this study, we propose P2U-SLAM, a framework for wide-FoV cameras that makes use of point and pose uncertainty to overcome weak feature correspondence introduced by a wide overlap area. The proposed system establishes a complete visual SLAM framework, which includes initialization, tracking, local mapping, loop closing, and map merging modules. To handle 3D points from the negative place (z<0z{<}0), the system inherits the Taylor camera model used in previous work. In addition, this work shows that it is mathematically justified and necessary to incorporate point and pose uncertainty during optimization. Moreover, the calculation and use of point and pose uncertainty are integrated into the SLAM framework, granting P2U-SLAM high positioning accuracy and robustness that exceeds the state-of-the-art methods on two public wide-FoV SLAM datasets. There is still a lot of room for improvement in P2U-SLAM. Despite a significant decrease in cumulative drift in the kilometer-level series, P2U-SLAM is still a pure visual SLAM framework. It is challenging for P2U-SLAM to work in some scenarios such as elevators and subways. Besides, the visualization of the map generated by P2U-SLAM is sparse and limited in downstream applications compared to dense maps.
In the future, we intend to extend P2U-SLAM to the visual-inertial framework to enhance the robustness of our algorithm in more extreme scenarios where vision fails. In addition, the uncertainty method of P2U-SLAM is strongly related to the current active research on 3D Gaussian Splatting, so achieving real-time dense mapping with higher efficiency from P2U-SLAM is also one of our exploration directions.